What is the pH of a mixture of 5 ml of 0.1 mol/litre sodium acetate and 4 ml of 0.1 mol/litre acetic acid? (pKa of acetic acid at 25oC = 4.76). How is the pH changed on adding 1 ml of 0.1 mol/litre HCl to the above mixture? (Show your answer by means of calculation)
1
Expert's answer
2019-08-16T05:41:44-0400

HCl→H+ + Cl-

"\frac {1L}{1000mL} \cdot \frac {0.1 mol}{1L} \cdot 1 mL = 0.0001 \space moles \space [H+]"


moles of CH3COOH = M*V = 0.1 mol/L * 0.004 L = 0.0004 moles

moles of CH3COO- = moles of CH3COONa = 0.1 mol/L * 0.005 L = 0.0005 moles

pKa=4.76

Ka=1.74*10-5





"Ka = \frac {[H^+] [CH_3COO^-]} {CH_3COOH]}"

"[H^+] = \frac {1.74 \cdot 10^ {-5} \cdot 0.0005}{0.0004} = 2.175 \cdot 10^ {-5}"

pH=-log10(2.175*10-5) = 4.66